OpenStudy (anonymous):

x(x^2-11y^2) simplify

OpenStudy (anonymous):

or u can use formula(a^2-b^2)=(a+b)(a-b)

OpenStudy (anonymous):

do I mulp2*2

OpenStudy (anonymous):

kepp x as it is but simplify (x^2-11y^2) using formula (a^2-b^2)=(a+b)(a-b)

OpenStudy (anonymous):

I do not know how to that

OpenStudy (anonymous):

as a^2-b^2)=(a+b)(a-b) x(x^2-11y^2)=x(x+11y)(x-11y) here a=x b=11y

OpenStudy (anonymous):

ok thanks is the end of it